OpenStudy (anonymous):

In the basketball game, Sarah made three times as many 2-point shots as 3-point shots. She made 8 shots in all. How many 2-point shots did she make? A. 2 B. 3 C. 5 D. 6 can someone explain to me how to work out these problems?

OpenStudy (jziggy):

Let's say that 2 point shots = x and 3 point shots = y Since she made 3 times as many 2 point shots as 3 points, x = 3y and since she made 8 shots in total, x + y = 8 we can substitute x = 3y in to the second equation to give 3y + y = 8 4y = 8 y = 2, So she made 2 of the 3 point shots. Substitute this answer into the second equation to give: x+2 = 8 x = 6, So she made 6 of the 2 point shots.